# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ 6 Applications and Numerical Experiments
In this section, we demonstrate the effectiveness of our proposed SDORE in the context of derivative estimation, nonparametric variable selection, and parameter identification in elliptic partial differential equations.
6.1 Derivative Estimates
@@
In this section we give a one-dimensional example, and a detailed example in two dimensions is shown in the supplementary material.
Example 6.1.
@@
Let the regression function be $f_{0}(x)=1+36x^{2}-59x^{3}+21x^{5}+0.5\cos(\pi x)$. The labeled data pairs are generated from $Y_{i}=f_{0}(X_{i})+\xi_{i}$ where $X_{i}$ are sampled from the uniform distribution on $[0,1]$, and $\xi_{i}$ are sampled from a normal distribution with the signal to noise ratio to be 30. The unlabeled data are also drawn from the uniform distribution on $[0,1]$. The regularization parameter is set as $\lambda=1.0\times 10^{-3}$.
@@
We conducted SDORE on this problem using 500 labeled data pairs and an additional 5000 unlabeled samples. The results are presented in Figure 1, which includes point-wise comparisons of function values and derivatives. Our SDORE method effectively estimates the regression function. In contrast, the errors in derivative estimates are more pronounced near the interval boundary. This is primarily due to the lack of observations of function values outside the intervals, preventing accurate estimation of the boundary derivatives. From a theoretical perspective, the convergence of the derivative in $L^{2}$-norm is guaranteed by Theorem 4.3. However, this theorem does not provide guarantees for accuracy on the boundary. Estimating the boundary error requires the interior estimation of second-order derivatives, as outlined in the trace theorem (Evans, 2010, Theorem 1 in Section 5.5).
@@
Figure 1: Numerical results of Example 6.1. (left) Noisy observations, regression function and SDORE. (right) Derivative function of the regression function and SDORE.
6.2 Nonparametric Variable Selection
@@
$f_{0}(x_{1},\ldots,x_{d})=f_{0}^{*}(x_{j_{1}},\ldots,x_{j_{d^{*}}}),\quad\{j_{1},\ldots,j_{d^{*}}\}\subseteq[d].$ (12)
This sparsity setting has garnered significant attention in the study of linear models and additive models, as extensively discussed in Wainwright (2019). In the context of reproducing kernel Hilbert space, Rosasco et al. (2010); Mosci et al. (2012); Rosasco et al. (2013) introduced a nonparametric variable selection algorithm. Nevertheless, their approach and analysis heavily depend on the finite dimensional explicit representation of the estimator, making it unsuitable for generalizing to deep neural network estimators.
We introduce the definition of relevant set, which was proposed by Rosasco et al. (2013, Definition 10). The goal of the variable selection is to estimate the relevant set.
Definition 6.1 (Relevant set).
@@
${\mathcal{I}}(f)=\Big{\{}k\in[d]:\|D_{k}f\|_{L^{2}(\nu_{X})}>0\Big{\}}.$
6.2.1 Convergence Rates and Selection Consistency
@@
Under Assumption 7, our focus is solely on estimating the low-dimensional function $f_{0}^{*}$ in (12) using deep neural networks. Consequently, the approximation and generalization error in Lemma 5.2 are reliant solely on the intrinsic dimension $d^{*}$. This implies an immediate result as follows.
Corollary 6.2.
Suppose Assumptions 1 to 7 hold. Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Assume that $f_{0}\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 2}$. Set the hypothesis class ${\mathcal{F}}$ as a ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,S)$ with $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d^{*}}{d^{*}+4s}})$. Let $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be the regularized empirical risk minimizer defined as (9) with regularization parameter $\lambda={\mathcal{O}}(n^{-\frac{s}{d^{*}+4s}})$. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}\Big{]}$ $\leq{\mathcal{O}}\Big{(}n^{-\frac{s}{d^{*}+4s}}\Big{)},$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}\Big{]}$ $\leq{\mathcal{O}}\Big{(}n^{-\frac{s}{2(d^{*}+4s)}}\Big{)}.$
Here we omit the logarithmic factors and the constant behind the big ${\mathcal{O}}$ notation is independent of $n$.
The convergence rate presented in Corollary 6.2 is solely determined by the intrinsic dimension $d^{*}$ and remains unaffected by the data dimension $d$, which effectively mitigates the curse of dimensionality when $d^{*}$ is significantly smaller than $d$.
@@
Corollary 6.3 (Selection Consistency).
Under the same conditions as Corollary 6.2. It follows that
@@
where the regularization parameter is given by $\lambda={\mathcal{O}}(n^{-\frac{s}{d^{*}+4s}})$.
Corollary 6.3 demonstrates that, given a sufficiently large number of data pairs, the estimated relevant set ${\mathcal{I}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})$ is equal to the ground truth relevant set ${\mathcal{I}}(f_{0})$ with high probability. In comparison, Rosasco et al. (2013, Theorem 11) only provided a one-side consistency
@@ 6.2.2 Numerical Experiments
In this section, we present a high-dimensional example which has sparsity structure to verify the performance of SDORE in variable selection. Throughout this section, let DORE denote the regressor where the Sobolev penalty is estimated only by the covariates in labeled data pairs (e.g. (2)).
Example 6.2.
@@
Let the regression function be $f_{0}(x)=\sum_{i=1}^{3}\sum_{j=i+1}^{4}x_{i}x_{j}$. Suppose the covariate in both labeled and unlabeled data are sampled from the uniform distribution on $[0,1]^{20}$. The label $Y$ is generated from the noisy model $Y=f_{0}(X)+\xi$, where $\xi$ is the noise term sampled from a normal distribution with the signal to noise ratio to be 25. The regularization parameter is set as $\lambda=1.0\times 10^{-2}$.
@@
We employed DORE and SDORE on this example, utilizing 100 labeled data pairs and an additional sample containing 1000 unlabeled covariate vectors. To assess the performance of these estimators, we evaluate DORE and SDORE on 100 distinct test sets, each comprising 100 test data pairs. Figure 2 visually presents the derivatives corresponding to each variable, revealing that the derivatives with respect to relevant variables $x_{1}$ to $x_{4}$ are significantly larger than those of the other variables. This shows that our proposed method can estimate the derivatives accurately, which in turn facilitates the variable selection. Additionally, Figure 2 displays the selection error, calculated as the mean of the false positive rate and false negative rate as defined by Rosasco et al. (2013), as well as the prediction errors in the form of root mean squared error (RMSE). Notably, the results consistently demonstrate the superior performance of SDORE over DORE, underscoring the advantages of incorporating unlabeled data.
@@
Figure 2: Numerical results of Example 6.2. (left) Derivatives estimation for the regression function $f_{0}$, which depends only on the $x_{1}$ to $x_{4}$. The dashed line is the threshold for variable selection. (center) Selection error on test set. (right) Prediction error on test set.
6.3 Inverse Source Problem
In this section, we aim at identifying the source density in elliptic problems from internal measurements. Inverse source problems have garnered significant attention in various scientific and engineering fields, as they arise in practical applications involving phenomena such as pollution source identification (Atmadja and Bagtzoglou, 2001; Badia and Ha-Duong, 2002; X. Liu, 2007; Barati Moghaddam, Mazaheri and Mohammad Vali Samani, 2021), dislocation problems (Badia and Hajj, 2013), and inverse problems of gravimetry (Isakov, 2018). Of particular relevance to this paper is the modeling of the seawater intrusion phenomenon (Badia et al., 2016; Costall et al., 2020), where the source term represents the pumping wells of freshwater within the context of seawater intrusion.
Consider the following second order elliptic equation with homogeneous Neumann boundary condition
$-\Delta u+w(x)u&=f(x),&&\text{in}~{}\Omega,\\ \nabla u\cdot\bm{n}&=0,&&\text{on}~{}\partial\Omega,\end{aligned}\right.$ (13)
where the potential function $w(x)$ is given. Further, the potential function $w(x)$ has a positive lower bound, that is, $w(x)\geq w_{0}>0$ for each $x\in\Omega$. Let $f_{*}$ be the ground truth source density and $u_{*}$ be the solution of (13) corresponding to the exact source $f_{*}$. The inverse source problems in elliptic equations aim to recover the unknown source density from noisy internal measurements ${\mathcal{D}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ generated from the following noisy model:
$Y_{i}=u_{*}(X_{i})+\xi_{i},\quad 1\leq i\leq n,$ (14)
where ${\mathbb{E}}[\xi_{i}|X_{i}]=0$ and ${\mathbb{E}}[\xi_{i}^{2}|X_{i}]<\infty$. Let $\mu$ denote the joint distribution of $(X_{i},Y_{i})$, and let $\mu_{X}$ denote the marginal distribution of $X_{i}$ for each $1\leq i\leq n$.
In this paper, we consider a variational approach proposed in Hu, Shu and Zou (2018). One first recovers the ground truth solution $u_{*}$ by the following optimization problem:
$\widehat{u}_{{\mathcal{D}}}^{\lambda}\in\operatorname*{arg\,min}_{u\in\operatorname{conv}({\mathcal{F}})}\frac{1}{n}\sum_{i=1}^{n}(u(X_{i})-Y_{i})^{2}+\lambda\|\nabla u\|_{L^{2}(\mu_{X})}^{2},$ (15)
which is equivalent to (11). Then a reconstruction $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ of the ground truth source density $f_{*}$ can be obtained by solving the following variational equation
$(\widehat{f}_{{\mathcal{D}}}^{\lambda},v)_{L^{2}(\mu_{X})}=(\nabla\widehat{u}_{{\mathcal{D}}}^{\lambda},\nabla v)_{L^{2}(\mu_{X})}+(w(x)\widehat{u}_{{\mathcal{D}}}^{\lambda},v)_{L^{2}(\mu_{X})},\quad\forall v\in H^{1}(\mu_{X}).$ (16)
In particular, according to Evans (2010, Section 6.1), if the reconstruction $\widehat{u}_{{\mathcal{D}}}^{\lambda}$ proposed in (15) satisfies $\widehat{u}_{{\mathcal{D}}}^{\lambda}\in H^{2}(\mu_{X})$, the variational equation (16) is equivalent to
$\widehat{f}_{{\mathcal{D}}}^{\lambda}=-\Delta\widehat{u}_{{\mathcal{D}}}^{\lambda}+w(x)\widehat{u}_{{\mathcal{D}}}^{\lambda}.$
In this work, the reconstruction $\widehat{u}_{{\mathcal{D}}}^{\lambda}$ is a deep neural network, and thus the condition $\widehat{u}_{{\mathcal{D}}}^{\lambda}\in H^{2}(\mu_{X})$ holds as long as the activation function has sufficient smoothness. Further, we can evaluate the derivatives of $\widehat{u}_{{\mathcal{D}}}^{\lambda}$ without error using the technique of automatic-differentiation (Rall, 1981; Griewank and Walther, 2008; Neidinger, 2010; Paszke et al., 2019).
6.3.1 Convergence Rates
We now turn to establish convergence rates for reconstructions $\widehat{u}_{{\mathcal{D}}}^{\lambda}$ and $\widehat{f}_{{\mathcal{D}}}^{\lambda}$. To begin, denote by $(H^{1}(\mu_{X}))^{*}$ the dual space to $H^{1}(\mu_{X})$. Then the dual norm of a function $u\in L^{2}(\mu_{X})$ is defined as
$\|u\|_{(H^{1}(\mu_{X}))^{*}}=\sup_{v\in H^{1}(\mu_{X})}\frac{(u,v)_{L^{2}(\mu_{X})}}{\|v\|_{H^{1}(\mu_{X})}}.$
The following corollary gives a convergence rate of the reconstruction adopted above.
Corollary 6.4.
Under Assumptions 1 and 2. Suppose Assumptions 3 and 6 hold for each function $u\in\{u_{*}\}\cup{\mathcal{F}}$. Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Assume that $u_{*}\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 2}$. Set the hypothesis class as a deep ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,S)$ with $L={\mathcal{O}}(\log n+\log(1+\sigma))$ and $S={\mathcal{O}}((1+\sigma^{-\frac{3d}{d+4s}})n^{\frac{d}{d+4s}})$. Let $\widehat{u}_{{\mathcal{D}}}^{\lambda}$ and $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be defined, respectively, as (15) and (16) with regularization parameter $\lambda={\mathcal{O}}((1+\sigma^{\frac{3s}{d+4s}})n^{-\frac{s}{d+4s}})$. Then it follows that
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{u}_{{\mathcal{D}}}^{\lambda}-u_{*}\|_{H^{1}(\mu_{X})}\Big{]}$ $\leq{\mathcal{O}}\Big{(}\Big{(}1+\sigma^{\frac{3s}{d+4s}}\Big{)}n^{-\frac{s}{d+4s}}\Big{)},$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{*}\|_{(H^{1}(\mu_{X}))^{*}}\Big{]}$ $\leq{\mathcal{O}}\Big{(}\Big{(}1+\sigma^{\frac{3s}{2(d+4s)}}\Big{)}n^{-\frac{s}{2(d+4s)}}\Big{)},$
where we omit the logarithmic factors and the constant behind the big ${\mathcal{O}}$ notation is independent of $n$ and $\sigma$.
Corollary 6.4 presents the stochastic convergence rates for the SDORE-induced reconstruction approach, providing explicit insights into the impact of the number of measurements $n$, and the variance proxy of noise $\sigma^{2}$. Inverse problems involving interior measurements with scattered data and random noise have garnered significant interest in the field of inverse problems (Chen, Tuo and Zhang, 2018; Chen, Zhang and Zou, 2022). However, the theoretical properties developed by Chen, Tuo and Zhang (2018); Chen, Zhang and Zou (2022) are specifically for finite element approximations (Brenner and Scott, 2008) in a fixed-design setting, and their analysis cannot be applied to neural network methods and the random-design setting.
6.3.2 Numerical Experiments
In this section, we present numerical reconstructions using method proposed in this paper. The accuracy of reconstructions are measured by the relative $L^{2}(\mu_{X})$-errors $\operatorname{err}(u)$, $\operatorname{err}(D_{k}u)$ and $\operatorname{err}(f)$ defined, respectively, as
$\operatorname{err}(u)=\frac{\|u-u_{*}\|_{L^{2}(\mu_{X})}}{\|u_{*}\|_{L^{2}(\mu_{X})}}\quad\text{and}\quad\operatorname{err}(\nabla u)=\frac{\|\nabla(u-u_{*})\|_{L^{2}(\mu_{X})}}{\|\nabla u_{*}\|_{L^{2}(\mu_{X})}}.$
Example 6.3.
We consider a two-dimensional problem, i.e., $d=2$. Let $\mu_{X}$ be the uniform distribution on $\Omega$. Set the ground truth solution $u_{*}(x_{1},x_{2})=\cos(2\pi x_{1})\cos(3\pi x_{2})$, and the potential $w\equiv 3\pi^{2}$. The noise term $\xi$ is drawn from Gaussian distribution $N(0,\sigma^{2})$ and is independent of $x=(x_{1},x_{2})$.
We utilize the SDORE-based approach to estimate the unknown solution $u_{*}$ from internal observations with noise levels of $\sigma=10\%$ and $\sigma=20\%$, respectively. A summary of the relative error is provided in Table 3. We observe that the error of the derivatives of SDORE is significantly smaller compared to the non-penalized estimator ($\lambda=0$), thereby demonstrating the effectiveness of the gradient penalty. Additionally, we find that the estimator with a large regularization parameter ($\lambda=1.0\times 10^{-2}$) fails to achieve accurate reconstruction. Moreover, the optimal regularization parameter for the case with $\sigma=20\%$ is larger than that for the case with $\sigma=10\%$. This suggests that larger regularization parameter is required when dealing with higher levels of data noise. These observations correspond well with our theoretical findings in Corollary 6.4.
Table 3: The relative $L^{2}$-error of reconstructions with various regularization parameters in Example 6.3.
Regularization parameters $\lambda$
$\sigma=10\%$ $1.0\times 10^{-2}$ $1.0\times 10^{-4}$ $1.0\times 10^{-6}$ $1.0\times 10^{-8}$ $0.0$
$\operatorname{err}(\widehat{u}_{{\mathcal{D}}}^{\lambda})$ $5.67\times 10^{-1}$ $3.61\times 10^{-2}$ $1.77\times 10^{-2}$ $1.79\times 10^{-2}$ $3.36\times 10^{-2}$
$\operatorname{err}(\nabla\widehat{u}_{{\mathcal{D}}}^{\lambda})$ $5.66\times 10^{-1}$ $7.84\times 10^{-2}$ $3.12\times 10^{-2}$ $\mathbf{3.10\times 10^{-2}}$ $8.97\times 10^{-2}$
$\sigma=20\%$
$\operatorname{err}(\widehat{u}_{{\mathcal{D}}}^{\lambda})$ $5.65\times 10^{-2}$ $3.90\times 10^{-2}$ $3.88\times 10^{-2}$ $3.48\times 10^{-2}$ $3.70\times 10^{-2}$
$\operatorname{err}(\nabla\widehat{u}_{{\mathcal{D}}}^{\lambda})$ $5.64\times 10^{-2}$ $\mathbf{4.73\times 10^{-2}}$ $5.03\times 10^{-2}$ $7.54\times 10^{-2}$ $5.10\times 10^{-2}$
7 Conclusion

### v2

@@ VI Applications and Numerical Experiments
In this section, we demonstrate the effectiveness of our proposed SDORE in the context of derivative estimation, and nonparametric variable selection.
VI-A Derivative Estimation
@@
In this section we give a one-dimensional example, and a detailed example in two dimensions is shown in Appendix F-A.
Example 1.
@@
Let the regression function be $f_{0}(x)=1+36x^{2}-59x^{3}+21x^{5}+0.5\cos(\pi x)$. The labeled data pairs are generated from a regression model $Y=f_{0}(X)+\xi$, where $X$ is sampled from the uniform distribution on $[0,1]$, and $\xi$ is sampled from a Gaussian distribution $N(0,\sigma^{2})$. Here the variance $\sigma^{2}$ is determined by a given signal-to-noise ratio $\frac{{\mathbb{E}}[f_{0}^{2}(X)]}{\sigma^{2}}=30$. The unlabeled data are also drawn from the uniform distribution on $[0,1]$. The regularization parameter is set as $\lambda=0.005$.
To demonstrate the effectiveness of SDORE in scenario where only few labeled sample is available, we conducted SDORE using 40 labeled data pairs and an additional 1000 unlabeled samples. The comparisons with the least-squares regression are presented in Figure 1, which includes point-wise comparisons of function values and derivatives. In the upper panel, the least-squares estimator generally matches the target function. However, the least-squares estimator fits the noise in the data rather than the underlying patterns near the left and right endpoints. Also, the lower panel shows that its estimated derivatives is inaccurate and unstable near the left and right endpoints. In comparison, our SDORE method successfully estimates the regression function and its derivatives simultaneously, and the regularization avoids the overfitting on the primitive function.
@@
The errors in derivative estimates by SDORE are more pronounced near the interval boundary. This is primarily due to the lack of observations of function values outside the intervals, preventing accurate estimation of the boundary derivatives. From a theoretical perspective, the convergence of the derivative in $L^{2}$-norm is guaranteed by Theorem IV.3. However, this theorem does not provide guarantees for accuracy on the boundary. Estimating the boundary error requires the interior estimation of second-order derivatives, as outlined in the trace theorem [72, Theorem 1 in Section 5.5].
@@
Figure 1: Numerical results of Example 1. (left) Scatter plot of noisy observations (paired data used for supervised learning), line plot of the ground-truth regression function and its values predicted by least-squares (LS) regression and SDORE. (right) The ground truth derivative function and its estimated values by LS and SDORE.
VI-B Nonparametric Variable Selection
@@
$f_{0}(x_{1},\ldots,x_{d})=f_{0}^{*}(x_{j_{1}},\ldots,x_{j_{d^{*}}}),\\ \{j_{1},\ldots,j_{d^{*}}\}\subseteq[d].$ (15)
This sparsity setting has garnered significant attention in the study of linear models and additive models, as extensively discussed in [91]. In the context of reproducing kernel Hilbert space, [23, 24, 25] introduced a nonparametric variable selection algorithm. Nevertheless, their approach and analysis heavily depend on the finite dimensional explicit representation of the estimator, making it unsuitable for generalizing to deep neural network estimators.
We introduce the definition of relevant set, which was proposed by [25, Definition 10]. The goal of the variable selection is to estimate the relevant set.
Definition 4 (Relevant set).
@@
${\mathcal{I}}(f)=\{k\in[d]:\|D_{k}f\|_{L^{2}(\nu_{X})}>0\}.$
VI-B1 Convergence Rates and Selection Consistency
@@
Under Assumption 7, our focus is solely on estimating the low-dimensional function $f_{0}^{*}$ in (15) using deep neural networks. Consequently, the approximation and generalization error in Lemma V.2 are reliant solely on the intrinsic dimension $d^{*}$. This implies an immediate result as follows.
Corollary VI.1.
Suppose Assumptions 1 to 7 hold. Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Assume that $f_{0}\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 2}$. Set the hypothesis class ${\mathcal{F}}$ as a ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,W,S)$ with $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d^{*}}{d^{*}+4s}})$. Let $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be the regularized empirical risk minimizer defined as (12) with regularization parameter $\lambda={\mathcal{O}}(n^{-\frac{s}{d^{*}+4s}}\log^{2}n)$. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}\Big{]}$ $\leq{\mathcal{O}}\Big{(}n^{-\frac{s}{d^{*}+4s}}\log^{4}n\Big{)},$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}\Big{]}$ $\leq{\mathcal{O}}\Big{(}n^{-\frac{s}{2(d^{*}+4s)}}\log^{2}n\Big{)}.$
The convergence rate presented in Corollary VI.1 is solely determined by the intrinsic dimension $d^{*}$ and remains unaffected by the data dimension $d$, which effectively mitigates the curse of dimensionality when $d^{*}$ is significantly smaller than $d$.
@@
Corollary VI.2 (Selection consistency).
Under the same conditions as Corollary VI.1. It follows that
@@
where $\lambda={\mathcal{O}}(n^{-\frac{s}{d^{*}+4s}}\log^{2}n)$.
Corollary VI.2 demonstrates that, given a sufficiently large number of data pairs, the estimated relevant set ${\mathcal{I}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})$ is equal to the ground truth relevant set ${\mathcal{I}}(f_{0})$ with high probability. In comparison, [25, Theorem 11] only provided a one-side consistency
@@ VI-B2 Numerical Experiments
In this section, we present a high-dimensional example which has sparsity structure to verify the performance of SDORE in variable selection. The additional experiments for variable selection are shown in Appendix F-B.
Example 2.
Let the regression function be
$f_{0}(x)=\sum_{i=1}^{3}\sum_{j=i+1}^{4}x_{i}x_{j}.$
@@
Suppose the covariate in both labeled and unlabeled data are sampled from the uniform distribution on $[0,1]^{20}$. The label $Y$ is generated from the regression model $Y=f_{0}(X)+\xi$, where $\xi$ is the noise term sampled from a Gaussian distribution with the signal-to-noise ratio to be 25, in the same way as Example 1. The regularization parameter is set as $\lambda=1.0\times 10^{-2}$.
In real-world applications, the process of labeling data can be prohibitively costly, resulting in a limited availability of labeled data. Conversely, there is an abundance of unlabeled data that is readily accessible. Hence, it becomes crucial to leverage few labeled data alongside a substantial amount of unlabeled data for the purpose of variable selection. Nevertheless, the task of variable selection with few labeled samples presents significant challenges. Due to the scarcity of data points, there is a restricted range of variability within the dataset, posing difficulties in accurately determining the variables that hold true significance in predicting the desired outcome.
To demonstrate the effectiveness of SDORE in this challenging scenario, we employ SDORE for the variable selection in this example, utilizing 50 labeled data pairs and an additional sample containing 100 unlabeled covariate vectors. Additionally, we use least-squares regression on the same data as a comparison. Figure 2 visually presents the empirical mean square (EMS) of estimated partial derivatives with respect to each variable on the test set, that is, for each $1\leq k\leq d$,
$\mathrm{EMS}_{k}=\frac{1}{n}\sum_{i=1}^{n}|D_{k}\widehat{f}(X_{i,k})|^{2}.$
@@
Here $\widehat{f}$ is an estimator, $\{X_{i}\}_{i=1}^{n}$ is a set of test data, and $X_{i,k}$ represents the $k$-th element of $X_{i}$. The results by SDORE reveals that the derivatives with respect to relevant variables $x_{1}$ to $x_{4}$ are significantly larger than those of the other variables, while least-squares regression wrongly regards $x_{9}$ as relevant variables, possibly due to the lack of paired training samples. This shows that our proposed method can estimate the derivatives accurately, which facilitates the variable selection. We select the variables by setting a 75% quantile threshold of the estimated partial derivatives. The partial derivatives greater than the threshold is considered relevant. Additionally, Figure 2 displays the mean selection error (SE), calculated as the mean of the false positive rate and false negative rate as defined by [25], as well as the root mean squared prediction error (PE) on the regression function. Notably, the results consistently demonstrate the superior performance of SDORE over least-squares regression, underscoring the advantages of incorporating unlabeled data.
Remark.
Since in Figure 2, for SDORE, the estimated partial derivatives with respect to the first four features is significantly larger than others, we can choose the threshold directly. In other application scenarios, if we can not observe such a clear difference, we can employ the strategy such as cross-validation (CV) to determine the number of features. Specifically, the cross-validation process involves dividing the dataset into a training set and a validation set independently. The training set is used for Sobolev regression, and the model’s performance is evaluated on the validation set. The mean square of partial derivatives, also known as the important score, is sorted from largest to smallest. The cross-validation process begins by selecting the feature with the largest important score, and adds the remaining most important features incrementally until the accuracy in the validation set no longer shows improvement.
@@
Figure 2: Numerical results of Example 2. The empirical mean square of the partial derivatives of the regression function $f_{0}$ (which depends only on the $x_{1}$ to $x_{4}$), estimated by least-squares fitting (LS, left) and SDORE (right). The dashed line is the 75% quantile threshold for variable selection. We also report the mean selection error (SE) for the estimated derivative function and the root mean squared prediction error (PE) for the primitive function by each method.
VII Conclusion